(1)
Solve each of the following systems by substitution
5x-2y=-5
y-5x=3




(2)
Solve each of the following systems by substitution
8x-4y=16
y=2x-4

Answer by HyperBrain(694)

(1)5x-2y=-5
(2)y-5x=3
(3)y=5x+3
Substituting (3) to (1),
5x+2(5x+3)=-5
5x+10x+6=-5
15x+6=-5
15x=-11
x=-11/15
y=5(-11/15)+3
=-2/3
The solution set, (x,y)=(-11/15,-2/3)


(1)8x-4y=16
(2)y=2x-4
Substituting (2) to (1),
8x-4(2x-4)=16
8x-8x+16=16
16=16
Therefore, the statement has infinite solutions!